OpenStudy (anonymous):

solve 2sqrtx-7 -10=-8 A.x=-6 B.x=-4 C. x=32 D.x=74

OpenStudy (anonymous):

Is it \[2\sqrt{x-7}-10=-8\]?

OpenStudy (anonymous):

yeaa

OpenStudy (anonymous):

\[2\sqrt{x-7}-10=-8\]\[2\sqrt{x-7}=2\]\[\sqrt{x-7}=1\]\[x-7=1\]\[x=8\]

OpenStudy (anonymous):

That's not one of my options that I have aboiv :/

OpenStudy (anonymous):

Are the numbers all right, is this the right equation?

OpenStudy (anonymous):

Yess

OpenStudy (anonymous):

I still get 8 :(

OpenStudy (anonymous):

:( that's weird